OpenStudy (anonymous):

The boys and girls at a middle school are raising money for technology at their school. The boys are selling cookie dough for $8 per box and the girls are selling candles for $5 each. They must raise more than $1000. The girls expect to sell at least 80 candles. Which of the following is a solution? A. 75 boxes of cookie dough and 90 candles B. 70 boxes of cookie dough and 85 candles C. 65 boxes of cookie dough and 80 candles D. 60 boxes of cookie dough and 75 candles

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

multiply 8 and 60 and 5 and 75 add it together and see if you get $1000

OpenStudy (anonymous):

got 855

OpenStudy (seaturtle113):

Nvm, sorry! I had very little sleep. :/ The answer would be A if you are asking what the solution would be that gets the kids the amount needed. because 75*8=600 and 90*5=450. 600+450=1050

OpenStudy (anonymous):

correct^